What is the mass of hydrochloric acid which reacts with 150 g of Zinc, which has a 95% Purity.
1
Expert's answer
2020-05-19T08:59:50-0400

Solution:

m(Zn) = 150 g × 0.95 = 142.5 g


The balanced chemical equation:

Zn + 2HCl = ZnCl2 + H2

According to the equation: n(Zn) = n(HCl)/2


Moles of Zn = Mass of Zn / Molar mass of Zn

n(Zn) = (142.5 g) / (65.38 g/mol) = 2.179 mol


n(HCl) = 2 × n(Zn) = 2 × 2.179 mol = 4.358 mol

Mass of HCl = Moles of HCl × Molar mass of HCl

The molar mass of HCl is 36.46 g/mol.

m(HCl) = (4.358 mol) × (36.46 g/mol) = 158.89 g = 158.9 g

m(HCl) = 158.9 g


Answer: 158.9 g is the mass of hydrochloric acid (HCl).